### v1

# Moving Object Detection from Moving Camera Using Focus of Expansion Likelihood and Segmentation

Abstract
Separating moving and static objects from a moving camera viewpoint is essential for 3D reconstruction, autonomous navigation, and scene understanding in robotics. Existing approaches often rely primarily on optical flow, which struggles to detect moving objects in complex, structured scenes involving camera motion. To address this limitation, we propose Focus of Expansion Likelihood and Segmentation (FoELS), a method based on the core idea of integrating both optical flow and texture information. FoELS computes the focus of expansion (FoE) from optical flow and derives an initial motion likelihood from the outliers of the FoE computation. This likelihood is then fused with a segmentation-based prior to estimate the final moving probability. The method effectively handles challenges including complex structured scenes, rotational camera motion, and parallel motion. Comprehensive evaluations on the DAVIS 2016 dataset and real-world traffic videos demonstrate its effectiveness and state-of-the-art performance.
I INTRODUCTION
Separating moving objects from static scenes in video is a fundamental task with applications in 3D reconstruction, obstacle avoidance for autonomous vehicle, and scene understanding for assistant robot. Previous methods [1] [2] [3] primarily rely only on optical flow information to differentiate object motion from camera motion. However, they often fail in complex, structured scenes, under intricate camera motion, or in low-textured environments. Because flow length depends on an object’s relative motion magnitude and distance from the camera, relying solely on flow makes it difficult to detect moving objects in complex 3D scenes. This work proposes a novel approach leveraging optical flow and segmentation to overcome these challenges. As shown in Fig. 1, the proposed method, Focus of Expansion Likelihood and Segmentation (FoELS), effectively detects moving objects in complex structured scenes.
Detecting moving objects in dynamic scenes is vital for various robotics applications, such as autonomous navigation and environmental understanding. While static scene segmentation has advanced significantly, identifying dynamic components remains challenging, particularly under complex conditions such as rotational motion, camera zoom, and cluttered backgrounds. The ability to accurately detect moving objects in dynamic scenarios facilitates precise reconstruction of the environment, which is invaluable for augmented and virtual reality applications.
To detect moving objects from a moving camera, it’s necessary to extract moiton in the image, and then remove camera-induced motion. Key challenges in this domain from the inherent complexity of real-world environments include:
  1. 1.
    Misinterpretation of large optical flow from nearby static objects: Large optical flow magnitudes from close, static objects can be erroneously interpreted as object motion.
  2. 2.
    Insufficient flow in low-textured regions: Environments with minimal texture hinder optical flow algorithms, leading to unreliable motion estimates.
  3. 3.
    Ambiguity in parallel motion: Objects moving parallel to the camera’s trajectory often produce optical flow that aligns with the background flow, causing detection ambiguities.
  4. 4.
    Detection of partially stationary objects: Objects with both moving and static parts (e.g., a walking animal with stationary limbs at certain moments) are challenging to classify accurately as moving, yet such distinction is crucial for applications like 3D reconstruction.
Figure 1: Sample result of FoELS. It detects moving objects from a moving camera at various distances within the scene.
II RELATED WORKS
This section reviews prior efforts in moving object detection. There are numerous methods for detecting moving objects using static cameras. For instance, Rozumnyi et al. [4] detect fast-moving objects in their research. However, there is limited research on moving object detection from a moving camera.
Notable review papers on moving object detection include [5], and [6]. Based on these reviews, we have identified several key areas of research in moving object detection, including flow orientation-based methods, focus of expansion (FoE) based approaches, and adversarial network methods.
Zhao et al. [6] categorized moving object detection application conditions into two types: detection of unseen scenes and detection of seen scenes. They mainly focused on the latter and background subtraction methods. However, they do not address how to handle moving backgrounds. MU-Net2 [7], one of the best approaches listed in their survey, is only effective for slightly moving cameras.
There is a similar task for moving object detection, namely “Semi-Supervised Video Object Segmentation on DAVIS”. The current best method for this task is HMMN [8]. However, this method requires human initialization, meaning it is not truly moving object detection but rather object tracking. Therefore, it falls outside the scope of this work.
While not evaluated in the aforementioned review papers, other notable approaches ([9, 1, 2, 10, 11, 3]) demonstrate notable effectiveness for moving object detection. They can be classified into the following three categories.
  1. 1.
    Flow orientation-based approach:
    Numerous methods utilize optical flow for moving object detection, such as [9]. Zhang et al. [1] introduced a technique that calculates optical flow orientation between adjacent video frames and reconstructs a background orientation field using Poisson fusion. This method aims to identify motion saliency by analyzing the discrepancies between the reconstructed background orientation and the observed orientation. While it works well for small camera movements, it fails when the camera moves straight ahead, where flow orientations are radially symmetrical.
  2. 2.
    FoE-based approach:
    FoE-based approaches estimate camera motion parameters such as rotation and translation. These methods assume a fixed FoE and identify moving objects by analyzing flow vectors relative to the FoE [2] [10]. Although conceptually robust, they struggle with scenarios involving unknown or dynamic FoE, limiting their utility in real-world conditions. While a direct FoE computation method (without optical flow) [11] was developed when optical flow was unreliable, current optical flow-based FoE estimation is more precise and prevalent, overcoming the former’s limitations (e.g., reliance on grayscale images and lack of quantitative evaluation).
  3. 3.
    Adversarial network approach:
    Yang et al. [3] leveraged adversarial learning frameworks to enhance motion detection. The generator-inpainter architecture trains the network to distinguish between moving and static regions by minimizing a loss function that encodes flow discrepancies. Despite achieving state-of-the-art performance on multiple existing datasets, it fails to detect moving objects in low-textured areas and generates false positives for close static objects when tested on our custom traffic video data.
All these representative methods rely primarily on optical flow for moving object detection, making them ineffective in complex, structured scenes. To address these limitations, we propose a novel approach that integrates both optical flow and segmentation information.
III PROPOSED METHOD
This section explains the system architecture and core algorithmic components.
III-A System Overview
Figure 2: System overview of the proposed method, FoELS.
Our system overview is illustrated in Fig. 2. The proposed pipeline consists of six main stages:
  1. 1.
    Optical Flow Estimation: Captures pixel-wise motion cues between consecutive frames.
  2. 2.
    Segmentation: Assigns class-specific prior moving probabilities and identifies static regions.
  3. 3.
    Camera Motion Detection: Determines if the camera is in motion by analyzing the optical flow ratio within static regions.
  4. 4.
    FoE computation Utilizes Random Sample Consensus (RANSAC) to compute the FoE from optical flow.
  5. 5.
    Moving Pixel Probability Estimation: An FoE-based moving pixel likelihood is computed from RANSAC outliers. This likelihood is then multiplied by segmentation-derived priors to yield the final moving pixel probability.
  6. 6.
    Object-Level Refinement: Validates moving pixel regions against panoptic segmentation results.
III-B Key Ideas to Overcome Challenges
In Section I, we identified four challenges for moving object detection from a moving camera. To address these challenges, FoELS integrates the following key ideas.
We incorporate an FoE-based approach, as it can address both straight-ahead camera motion and the misinterpretation of large optical flow from nearby static objects (Challenge 1). However, unlike [10], we do not assume a fixed FoE and instead allow it to vary with each frame. This allows us to handle various camera movements. To address the challenges of misinterpreting large optical flow from nearby static objects (Challenge 1) and insufficient flow in low-textured regions (Challenge 2), which [3] struggled with, we incorporate segmentation as an additional cue for detecting moving objects. Additionally, we introduce a method to address the challenge of ambiguity in parallel motion (Challenge 3), which will be explained in III-H. Finally, we add object-level refinement, which extracts moving objects after identifying moving pixels to address the challenge of detecting partially stationary objects (Challenge 4). The method will be explained in III-I. This paper’s primary objective is the detection of complete moving objects, rather than individual moving pixels, t o support downstream applications such as 3D reconstruction.
Ultimately, FoELS integrates optical flow and segmentation to overcome the limitations of previous methods.
III-C System Details
First, frames $t-1$ and $t$ are used to compute the optical flow. Simultaneously, segmentation is performed on frame $t$, and each pixel is assigned a prior moving probability according to a manually predefined class-moving probability table. Sky regions identified through segmentation are removed, since optical flow cannot be computed there. Concurrently, static areas (e.g., ground, mountains, and buildings) are identified, and the flow within these regions is analyzed. If the flow existing ratio in the static area exceeds a specified threshold, the camera is considered to be in motion, and the FoE is computed using RANSAC. The inliers from the RANSAC process are attributed to camera motion, while the outliers are considered as moving pixels. Once the moving pixels are identified, we map them back to moving objects using the panoptic segmentation results.
The detailed flow chart of the above procedure is listed in Fig. 3.
Figure 3: Detailed flowchart of the proposed method, with color coding corresponding to Fig. 2. The side image illustrates the process outlined in the flowchart.
III-D Optical Flow Estimation
Accurate optical flow estimation forms the foundation of FoELS. We adopt the UniMatch model [12], which is trained on challenging benchmarks such as Sintel and KITTI, and achieves state-of-the-art accuracy. UniMatch provides dense flow maps, capturing subtle motion patterns critical for subsequent analysis.
III-E Segmentation
Segmentation is a widely applicable and extensively investigated area in computer vision. Numerous studies have focused on segmentation [13, 14, 15, 16, 17, 18, 19]. Common semantic segmentation approaches proved insufficient for detecting moving objects, as it is crucial to distinguish between individual object instances, some of which may be moving while others remain static. Therefore, panoptic segmentation, which combines semantic and instance segmentation, was adopted for this work.
Several segmentation models capable of handling panoptic segmentation tasks have been proposed recently, such as Mask2Former [20]. Among these advanced models, OneFormer [21] is currently the state-of-the-art panoptic segmentation model. Consequently, the OneFormer panoptic segmentation model is utilized in this approach. Each class is assigned a prior probability reflecting its tendency to be dynamic. For instance, sky and building classes have low moving probabilities, while vehicle and pedestrian classes have higher values.
In FoELS, these class-based prior moving probabilities are manually defined. These values were determined and adjusted based on several experiments, and the same predefined values are used for all datasets.
III-F Camera Motion Detection
After obtaining segmentation results and assigning prior moving probabilities, static areas are identified as regions where the moving probability is below a manually defined threshold. The camera is considered to be moving if the ratio of existing optical flow in these static areas exceeds a manually defined threshold. This work does not compute the camera’s egomotion; instead, it only determines whether the camera is moving. This determination is sufficient for computing the FoE and identifying moving objects.
III-G FoE Computation
To compute the FoE, two optical flow vectors within the identified static area are selected. An initial FoE candidate is determined as the intersection point of the lines extended from these two flow vectors. The sign of this FoE candidate (positive for a source of optical flow, negative for a sink) is concurrently determined from the directions of these flow vectors. Without this sign, objects moving in the opposite direction cannot be correctly identified as moving. Finally, the RANSAC algorithm is employed to robustly estimate the FoE from the set of available flow vectors in the static regions.
III-H Moving Pixel Probability Estimation
In this work, we employ the terms “prior”, “likelihood”, and “posterior” in a manner analogous to Bayes’ theorem. Specifically, the segmentation-based moving probability is defined as the “prior”, and the FoE-based moving probability is defined as the “likelihood”. Their product is subsequently termed the “posterior moving pixel probability” (or simply “moving pixel probability”). It is important to note that while this nomenclature is adopted due to the architectural resemblance of our approach to Bayesian updates, FoELS is not a strict Bayesian inference model. This distinction arises because the “likelihood” in our framework is not conditioned on the “prior”.
The moving likelihood, computed from the FoE, is multiplied by segmentation-derived priors to yield posterior moving pixel probabilities. This FoE-based moving likelihood is determined from outliers identified during the RANSAC FoE computation. These outliers correspond to points where the observed optical flow angle deviates from that expected under the computed FoE.
However, relying solely on this angular difference is insufficient for accurately handling scenarios involving motion parallel to the camera, where the flow direction of the moving object closely aligns with that of the background. Fig. 10 (d) exemplifies such a problematic scene. Though the detail explanation of the figure is in section IV-C, see the lower part of the truck in the center image, which shows the FoE inlier and outliers. Where the optical flow vectors (green arrows indicating FoE inliers) for portions of the truck align with the background flow, signifying identical flow angles. To address this ambiguity, information regarding differences in flow length is incorporated. Directly multiplying probabilities derived from length differences can induce false positives for nearby static objects, as their flow magnitudes are often large. Therefore, a logarithmic factor of the length difference is added to the angle-based moving likelihood, and the sum is subsequently clipped to the range [0, 1]. This approach, as demonstrated in Fig. 10 (d), enables FoELS to successfully detect the truck as a moving object. The method primarily emphasizes the flow angle difference while also considering significant flow length discrepancies, particularly for detecting parallel motion.
The moving pixel probability is computed as follows:
$P_{M}$ $\propto P_{seg}\cdot P_{FoE},$ (1)
$P_{FoE}$ $=\text{clip}_{[0,1]}(P_{a}+\alpha F_{l}),$ (2)
where, $P_{M}$ is the (posterior) moving pixel probability. $P_{seg}$ denotes the segmentation-based prior probability, and $P_{FoE}$ is the moving pixel likelihood based on the FoE. The function $\text{clip}_{[0,1]}()$ clips a value to the range [0, 1]. $P_{a}$ is the angle-based probability, which will be detailed below. $\alpha$ serves as the weighting factor for the flow length, and $F_{l}$ is the length factor, also detailed below.
Here, the angle-based probability $P_{a}$ is calculated proportionally to the optical flow angle difference between the optical flow at the point and the expected flow direction based on the FoE. $P_{a}$ is normalized to become 0.5 at a predefined angle difference threshold, $\theta_{th}$. The length factor, $F_{l}$, incorporates the base-10 logarithm of the relative flow length. This logarithmic scaling allows for the consideration of significant flow magnitude differences, pertinent for parallel motion, while diminishing the influence of minor variations. These components are specifically formulated as:
$P_{a}$ $=\text{clip}_{[0,1]}(0.5\cdot d_{a}/\theta_{th}),$ (3)
$F_{l}$ $=|\text{log}_{10}(|d_{l}|)|,$ (4)
$d_{a}$ $=\arccos\left(\frac{\mathbf{v}_{F}\cdot\mathbf{v}_{P}}{||\mathbf{v}_{F}||\cdot||\mathbf{v}_{P}||}\right),$ (5)
$d_{l}$ $=||\mathbf{v}_{P}||/\overline{||\mathbf{v}_{P,static}||},$ (6)
where, $\mathbf{v}_{F}$ is the vector from the FoE to the point, and $\mathbf{v}_{P}$ is the optical flow vector at the point. The term $d_{a}$ represents the angular difference calculated from these vectors. $d_{l}$ is the relative flow length difference, by $\overline{||\mathbf{v}_{P,static}||}$, which denotes the mean optical flow magnitude observed in static regions.
All thresholds were empirically determined. In our experiments, the weighting factor $\alpha$ was set to 0.25, and the angle threshold $\theta_{th}$ to 30 degrees.
III-I Object-Level Refinement
Finally, the computed moving pixel probabilities are aggregated to an object level. This step is crucial for ensuring that an entire object is classified as moving, even if only a portion of it exhibits detectable motion (addressing Challenge 4). To achieve this, a binary moving pixel mask is first generated by thresholding the posterior moving pixel probability $P_{M}^{\prime}$. A threshold of $0.250.5^{2}=$ 0.25 is used for $P_{M}^{\prime}$, reflecting the fact that $P_{M}^{\prime}$ is a product of two probabilities ($P_{seg}$ and $P_{FoE}$); this threshold implies that both contributing probabilities are at least 0.5. Subsequently, an object-level moving mask is derived. For each object instance identified by the panoptic segmentation, the percentage of pixels within that instance that are marked as moving in the binary pixel mask is calculated. If this percentage exceeds a threshold of 0.01, the entire object instance is classified as moving. This low threshold is employed to effectively detect objects where only a small part is in motion, such as the tail of an animal or a limb of a person, which can sometimes constitute as little as approximately 3% of the total object area.
III-J Comparison of Tractable Scenes
To provide a concise comparison of the advantages and disadvantages of related works and the proposed method based on tractable scenes, we present a comparison table of tractable scenes in TABLE I. The proposed method, FoELS, is capable of handling a broader range of scenarios compared to existing methods.
TABLE I: Comparison of tractable scenes. $\times$: Not tractable, $\triangle$: Partially tractable, ✓: Tractable
The possible reasons for tractability are listed in the bottom row for FoELS.
Method Stop Go Forward Rotate Go Forward and Rotate Textureless object Close object Close dominant object
Flow Orientation [1] $\times$ $\times$ $\times$ $\times$ $\times$
FoE [10] $\times$ $\times$ $\times$ $\times$
AdversarialNet [3] $\triangle$ $\triangle$ $\times$ $\times$ $\times$
FoELS (Ours) $\triangle$ $\triangle$ $\times$
by Orientation by FoE by Seg by FoE
IV EVALUATION
IV-A Datasets
Experiments were conducted on the DAVIS 2016 dataset [22], the FBMS-59 dataset [23], and a custom-collected traffic video dataset. The DAVIS 2016 and FBMS-59 datasets, which are annotated for moving objects, were utilized for quantitative evaluation. These relatively small datasets pose a risk of overfitting for training-based approaches. Though our method involves fitting only a few parameters, rather than comprehensive training, this risk is pertinent to the training-dependent methods against which we compare. To evaluate robustness and applicability in real-world scenarios, the custom traffic video dataset, which is unannotated, was used for qualitative assessment.
IV-A1 Quantitative Evaluation Dataset
For quantitative evaluation, we utilized the DAVIS 2016 dataset and the FBMS-59 dataset.
The FBMS-59 dataset provides annotations specifically for the moving object detection task. In contrast, the DAVIS 2016 dataset is primarily designed for video object segmentation, which is a binary labeling problem focused on separating foreground object from the background in a video. Consequently, the foreground annotations in DAVIS 2016 may sometimes include objects that are part of a moving background.
Upon careful examination of the DAVIS 2016 dataset, it was observed that certain scenes are inappropriate for evaluating moving object detection due to the presence of unannotated moving backgrounds. For instance, the breakdance scene features background spectators in motion who are not labeled as moving objects. The dataset comprises 50 scenes in total. After identifying and excluding scenes with significant unannotated background motion, the following three scenes were removed: bmx-bumps, breakdance, and dance-jump (3 out of 50).
Furthermore, an additional 15 scenes exhibit slight, unannotated background motion. However the background movements in these scenes are minor, and to maintain a substantial dataset size for evaluation, they were retained in our evaluation set. Consequently, the final evaluation set, termed DAVIS 2016 train-val-movobj, consists of the remaining 47 scenes.
IV-A2 Qualitative Evaluation Dataset
To assess the performance of FoELS in real-world conditions, a custom traffic video dataset was captured. This dataset encompasses a variety of challenging scenarios, including vehicles moving parallel to the camera, stopped vehicles, and operation in low-textured environments. Furthermore, it uniquely features instances of camera zooming, a condition not typically represented in standard moving object detection datasets such as DAVIS 2016 and FBMS-59.
IV-B Evaluation Metrics
We adopt Intersection-over-Union (IoU) scores as the primary evaluation metric, consistent with the methodology employed by the Adversarial Network [3]. This facilitates a direct comparison of FoELS’s performance against that of the Adversarial Network. The scene IoU score is calculated by averaging the IoU scores across all frames within a sequence. The final IoU score is subsequently determined by averaging all computed scene IoU scores.
IV-C Results
TABLE II: Quantitative evaluation result. The values represent the average IoU scores over the DAVIS 2016 train-val-movobj sequences and FBMS-59 Testset scenes.
DAVIS 2016 FBMS 59
Adversarial Net 0.599 0.369
FoELS (Ours) 0.757 0.695
The final quantitative evaluation results are presented in TABLE II. The Adversarial Network’s training protocol included the use of test data. In contrast, FoELS was trained without access to test data and employed consistent settings across all datasets. Despite this difference in training methodology, FoELS surpassed the state-of-the-art Adversarial Network method, achieving a higher IoU score.
Fig. 4 shows an example of the visual results from the above evaluation. This figure illustrates the step-by-step results of the process detailed in Section III. The first row displays: (a) the input frame, (b) the segmentation result, where different colors denote distinct classes, and (c) the prior moving probability derived from segmentation. The prior probability is visualized using a jet colormap, where red indicates higher probability and blue signifies lower probability. The second row presents: (d) the optical flow, with orientation encoded by color, (e) the optical flow field highlighting Focus of Expansion (FoE) inliers (green arrows) and outliers (red arrows). An existing FoE in the image is marked with a thick red cross. (f) The FoE-based moving likelihood, also depicted using a jet colormap. The third row shows: (g) the posterior moving pixel probability, calculated as the product of the prior moving probability and the FoE-based moving likelihood, (h) the refined object-level moving mask, demonstrating the aggregation of moving pixels to an object level, and (i) the final moving object mask overlaid on the input image. In this particular example, the bear’s hand remains stationary while the bear is walking. Nevertheless, FoELS successfully extracts the entire bear due to the object-level refinement process.
Figure 4: Example visual results of FoELS on the DAVIS 2016 bear scene. First row (left to right): (a) Input frame, (b) segmentation result, and (c) prior moving probability derived from segmentation. Second row (left to right): (d) Optical flow, (e) optical flow with FoE inlier (green arrows) and outliers (red arrows), and (f) the FoE-based moving likelihood. Third row (left to right): (g) Posterior moving pixel probability, (h) refined object-level moving mask, and (i) the final moving object result.
Fig. 5 compares the results of the Adversarial Network with those of FoELS for the same scene as Fig. 1. The left side shows the results of the Adversarial Network, while the right side displays the results of FoELS. The Adversarial Network falsely detects nearby vegetation and poles as moving objects due to their significantly different optical flow compared to the background. In contrast, FoELS successfully identifies only the genuinely moving objects by primarily relying on FoE-based flow orientation analysis.
Figure 5: Comparison results with Adversarialnet (left) and FoELS (right).
Fig. 6 shows the step by step visualization results of the same scene as Fig. 1. In this example, it can be seen why FoELS can correctly detect cars almost moving parallel to the camera, and the nearby static pole, despite exhibiting large optical flow, is correctly identified as stationary.
Figure 6: Example visual result of FoELS on custom traffic video. Each image corresponds to the same visualization format as shown in Fig. 4.
Fig. 7 compares the results of the Adversarial Network with those of FoELS on the custom zoom in/out video evaluation, where the camera remains stationary while zooming. The left panel illustrates the results of the Adversarial Network, while the right panel displays the results of FoELS. In this initial frame, the camera is nearly stationary, and no zoom is applied. However, the Adversarial Network produces numerous false positives. This indicates a lack of robustness in the Adversarial Network when applied to novel, untrained scenes. Conversely, FoELS exhibits no false positives in this scenario.
Figure 7: Comparison results: Adversarial Network (left) and FoELS (right) on a custom zoom in/out video (initial frame).
Fig. 8 presents a similar comparison between the Adversarial Network and FoELS. In this instance, the camera is actively zooming in while the train is in motion. An incoming train is positioned near the center of the image, while simultaneously the background exhibits motion due to the camera zoom. Notably, the Adversarial Network fails to detect any moving objects. In contrast, FoELS successfully identifies the approaching train while correctly disregarding the background motion induced by the zoom.
Figure 8: Comparison results: Adversarial Network (left) and FoELS (right) on a custom zoom in/out video (train approaching during zoom).
Fig. 9 illustrates the intermediate processing steps for the same frame of Fig. 8, employing the visualization format detailed in Fig. 4. This visualization demonstrates the successful detection of the moving train by FoELS.
Figure 9: Zoom in/out scene visual result of FoELS on the same frame with Fig. 8. Each image corresponds to the same visualization format as shown in Fig. 4.
We present additional visual results of FoELS in Fig. 10. Fig. 10 (a) shows an example of the black swan scene in the DAVIS 2016 dataset. The swan’s color is very close to the background river, therefore hard to segment the swan. Prior to selecting the final segmentation model, we evaluated several state-of-the-art approaches and found that OneFormer [21], the model ultimately adopted, successfully segments the swan, thereby enabling FoELS to detect its motion even in this challenging scene.
Fig. 10 (b) illustrates an example of the camel scene in the DAVIS 2016 dataset. In this scene, it successfully detects only the moving camel while ignoring the static camel.
Other scenes also contain challenging scenarios, where some potential moving objects are in motion while others remain static. However, FoELS successfully detects the moving objects in all of them.
((a)) DAVIS:blackswan
((b)) DAVIS:camel
((c)) DAVIS:dance-twiri
((d)) Custom:truck
((e)) FBMS:cars4
((f)) FBMS:tennis
Figure 10: Example visual results of FoELS. Each image corresponds to the same visualization format as shown in Fig. 4.
IV-D Ablation Studies
We conducted ablation studies to evaluate the effectiveness of each component of FoELS. The results are presented in TABLE III. The study began with a comparison of semantic and panoptic segmentation. Specifically, we evaluated the semantic segmentation models InternImageT, as well as the panoptic segmentation model OneFormer. For the OneFormer model, object refinement was subsequently incorporated. Finally, the inclusion of the FoE sign and further parameter adjustments constituted the final FoELS configuration.
TABLE III: Ablation study results. The values represent the average IoU scores over the DAVIS 2016 train-val-movobj sequences. ”OneFormer with ObjRefine” refers to the OneFormer model for panoptic segmentation with object refinement. ”+ FoE sign” indicates the addition of the FoE sign, representing the final FoELS configuration.
IoU
InternImageT (Semantic) 0.532
Oneformer (Panoptic) with ObjRefine 0.65
+ FoE sign (=FoELS) 0.757
V CONCLUSION
FoELS presents an innovative method for detecting moving objects from a moving camera, seamlessly integrating optical flow, segmentation, and camera motion detection through FoE estimation. By addressing challenges such as rotational motion and low-textured environments, FoELS demonstrates robust performance across diverse scenarios. Owing to its FoE-centered flow analysis, FoELS can detect objects even during camera zoom operations, a scenario often challenging for existing moving object detection techniques. FoELS demonstrates robust performance on the DAVIS 2016 and FBMS-59 datasets, as well as real-world traffic videos, employing consistent settings across all datasets, underscoring its potential for various applications in robotics and computer vision. Furthermore, the modular architecture of FoELS, which is not tightly coupled with specific segmentation or optical flow methods, allows for the integration of future advancements in these areas, potentially leading to further performance enhancements.
Future work will focus on optimizing FoELS for real-time performance, enhancing camera motion detection (e.g., via direct estimation using deep neural networks as an alternative to FoE-based inference), and integrating object tracking (e.g. [24, 25]) to leverage temporal information for improved accuracy and consistency.